2) A piece of plastic has a net charge of +2.3 μC caused by an deficit of electrons. How many more protons than electrons does this piece of plastic have?
Answer #1

the charge of 1 electron = 1.6*10^-19 C

so the number of deficient electrons = 2.3*10^-6 / 1.6*10^-19

= 1.437*10^13